OpenStudy (anonymous):

PLEASE HELP WITH THIS VECTORS PROBLEM! Explain how you would solve the following problem using rectangular vector components. Your friend swims diagonally across a 25m by 10m pool at 1m/s. How fast would you have to walk around the edges of the pool to get to the same point at the same time as your friend. Please show all work, thanks :)

OpenStudy (turingtest):

just break the velocity into components as we did last time

OpenStudy (anonymous):

25/1 and 10/1?

OpenStudy (anonymous):

First, find the length of the diagonal and the length of the semiperimeter. Those are the two distances. You can solve for the time which is the same for both, then solve for the other velocity.

OpenStudy (turingtest):

you just need to think about the angle that he is swimming at

OpenStudy (anonymous):

so pythagorean theorem?

OpenStudy (anonymous):

Hmm, well, if you have to solve it using components (which is needlessly complicated, but good practice) then you might have to break out the sines and cosines..

OpenStudy (anonymous):

Please just show me the fastest way

OpenStudy (anonymous):

But, yeah, the easy way is to find the diagonal using P.T. and treat it as a simple d=r*t problem, but from the wording, I think it wants you to do it the hard way.

OpenStudy (anonymous):

Ive drawn this so far, 

OpenStudy (anonymous):

Fastest ways is what I said at first: Find the two distances, find the time, then the only unknown in the equation is the second velocity.

OpenStudy (anonymous):

Good, draw in the diagonal and use PT to find its length.

OpenStudy (anonymous):

So 25^2+10^2=sqrt725=26.93

OpenStudy (anonymous):

Right - best to leave it in radical form for now so you don't introduce round-off error too soon. The swimmer's velocity is 1, so the time to cross in seconds is the same as the distance. Now find the walker's distance, divide by that amount of time and that is the walker's velocity.

OpenStudy (anonymous):

That'll get you a quick answer, but it doesn't really follow the directions given which state to use vector components.. If you'll get points deducted on this assignment for not following directions, then you might have to go ahead and do it the hard way anyway - but at least you'll know what the right answer is ahead of time so you can check your work.

OpenStudy (anonymous):

Ok, can we do it the way the question asks then

OpenStudy (turingtest):

I assumed you were allowed to have two different velocities around the pool, so I'm not sure I understand the problem.

OpenStudy (anonymous):

It asks for the velocity of the walker

OpenStudy (turingtest):

one velocity for the horizontal portion, one for the vertical

OpenStudy (turingtest):

if it is supposed to be just a single velocity the whole trip I have to rethink it...

OpenStudy (anonymous):

wouldnt that just be the avg velocity?

OpenStudy (anonymous):

Ultimately, I think you can state the answer as a single constant velocity, but I see no reason why the walker can't change speed after rounding the corner..

OpenStudy (anonymous):

You could give the walker vector components of velocity that match the swimmer.. (again, seems needlessly complicated for this problem) \[v_{swimmer}=(1m/s)(25/\sqrt{725}_{horizontal}+10/\sqrt{725}_{vertical})\]

OpenStudy (anonymous):

Ok, for that, I got: 1.3